# Vaidya-Reissner-Nordström Extension On the White-hole Region and the White Hole Evaporation Dynamic

### v1

@@ III.2 Mass Loss and Outgoing Energy Flux
In the outgoing Vaidya picture, $M(u)$ encapsulates the evaporation law. For example, one might prescribe a linear decrease $M(u)=M_{0}-\alpha u$ during some interval or an asymptotic decay $M(u)\sim M_{0}-\beta\ln(1+u)$, etc. The rate $|M^{\prime}(u)|$ is related to the power carried by the radiation. In particular, far from the hole, the luminosity measured at infinity is $-\frac{dM}{du}$ (since $M^{\prime}(u)<0$). This matches the idea that the mass-energy radiated to infinity reduces the Bondi mass at future null infinity. In the stress tensor above, $T_{uu}=-\frac{M^{\prime}(u)}{4\pi r^{2}}$ is actually an energy flux density along the outgoing null direction $u$.
Then we review Hiscock’s evaporating hole model. W. A. Hiscock in 1981 constructed [11] explicit piecewise Vaidya solutions to model an evaporating black hole. In Hiscock’s Model, the spacetime is divided into regions: an initial static Schwarzschild region (mass $M_{0}$), a Vaidya region where $M(u)$ decreases (modeling Hawking radiation emission), and a final flat spacetime region after the mass has gone to zero. This provided a simplified picture of a black hole that forms (or exists) and then evaporates completely in a finite time. Notably, these classical solutions have event horizons of finite duration – unlike an eternal horizon, the horizon “turns off” once the mass vanishes. The outgoing radiation is carried by null dust stress-energy, as in the expressions above. Hiscock’s analysis showed some important consequences of different evaporation scenarios: if one naively lets $M(u)$ reach zero at a finite retarded time $u_{\text{end}}$, the endpoint of evaporation is marked by a pathological surge of energy. In fact, studying quantum fields on this background reveals an infinite flux of particles created along the Cauchy horizon (the would-be last light ray from the evaporation) if $M$ vanishes abruptly at finite $u$. This suggests that a realistic evaporation might occur more gradually (e.g. $M\to 0$ only as $u\to+\infty$).
@@
At large $r$, $\theta_{(n)}\approx-1/r<0$ which ingoing rays converge as in flat space. But if we approach the radius $r=2M(u)$, the ingoing expansion $\theta_{(n)}$ approaches zero and changes sign. Inside the white hole, one actually has $\theta_{(n)}>0$ (an anti-trapped region where even ingoing light is forced outward). The surface $r=2M(u)$ thus acts as a marginal surface. In fact, $r=2M(u)$ is a dynamical apparent horizon for the spacetime: it is the boundary between normal space and the anti-trapped region. For the ingoing Vaidya black hole case, $r=2M(v)$ is a marginally trapped surface. In either case, the Vaidya horizon satisfies conditions like $\theta_{(\ell)}\theta_{(n)}=0$ on it. We identify $r_{h}(u)=2,M(u)$ as the apparent horizon radius at a given retarded time $u$. Differentiating this:
@@ III.3 Horizon Dynamics and Spacetime Structure during Evaporation
In a dynamic spacetime, the event horizon is defined globally, whereas an apparent horizon or trapping horizon is defined locally. In the Vaidya solution, a natural apparent/anti-trapping horizon is given by $r=2M(u)$ as noted. For an evaporating white hole, this is could be defined as the past apparent horizon. In an idealized evaporation that leaves no remnant, the event horizon coincides with this surface initially but will end at some point. Specifically, if $M(u)$ goes to zero at $u_{\text{end}}$, then beyond that point there are no trapped or anti-trapped surfaces, means the spacetime becomes Minkowski with no horizon. Thus the event horizon terminates at a finite advanced time. Even if $M(u)$ only asymptotically approaches 0 as $u\to\infty$, the horizon effectively shrinks to zero size in the infinite future.
@@
In summary, for evaporating black hole, horizon is found to be timelike in the Vaidya solution (to first order in $M^{\prime}$). This means an outside observer could actually cross the retreating horizon if they move fast enough outward, since the horizon is moving inward. Intuitively, the horizon is contracting, so outgoing light just outside can outrun it. A timelike horizon also implies it has a surface area that decreases with time, when $dA<0$ we say it as timelike membrane. This is a violation of the classical area theorem, but it is possible here because the null energy condition is violated by Hawking-like flux.
@@
For shrinking white hole, horizon is spacelike in analyses of dynamic solutions. A spacelike horizon recedes superluminally to the past. In the time-forward picture, it means the horizon’s boundary moves at faster-than-light rate inward. In essence, the white hole’s anti-trapping horizon vanishes so quickly that it lies within the backward light cone of its boundary. Physically, once the white hole has emitted most of its mass, the remaining horizon collapses “faster than light” in the diagram, so no world line can remain on the horizon, as it is a fleeting, space-like boundary. This corresponds to a rapid evaporation or an endpoint of the white hole. By contrast, an expanding white hole, which would have increasing $M$ if such were considered, yields a timelike horizon, analogous to an evaporating Black hole’s horizon being timelike.
@@
These properties were confirmed in recent studies of Kerr–Vaidya spacetime, which reduce to spherical symmetry in the limit $a\to 0$. There, it was found that a shrinking white hole’s horizon is spacelike, whereas a slowly expanding one is timelike. Similar, a slowly evaporating black hole’s apparent horizon is timelike. All these dynamic horizons differ from the static case, which the classical event horizon will wrinkles and tilts due to the mass flux.
III.4 Penrose diagram and global structure
@@
The spacetime of an evaporating white hole can be visualized as follows. Start with a past singularity (much like the white hole singularity in the past of an eternal Schwarzschild). Emanating from this $r=0$ singularity is the white hole event horizon (a past horizon) which initially might be near $r=2M_{\text{initial}}$. As time progresses, the horizon moves outward for a while (if the white hole initially increases to $M_{\text{initial}}$) and then begins to shrink as $M$ decreases. Outside the horizon, there is an outflow of radiation to infinity. If the evaporation is complete, at some retarded time $u=u_{+}$ the mass approaches zero. At this point, the horizon meets the center $r=0$. In a Penrose diagram, the horizon might terminate at a point on future null infinity or at future timelike infinity depending on how $M$ vanishes. In any case, after evaporation, the spacetime becomes flat (or at least free of any horizon or singularity) – effectively connecting to a region isometric to a portion of Minkowski space. No future singularity remains, since the original singularity was in the past. In this sense, the white hole has completely disappeared, leaving behind empty space (plus outgoing radiation that has dispersed).
@@ IV Vaidya-Reissner-Nordström white hole
The nonzero components of the Einstein tensor $G_{\mu\nu}$ for the above metric can be computed, as we assuming $r(u,v)$ and $f(u,v)$ are smooth. Using $g_{uv}=-f$ and set $18\pi G=c=$ 1 units for brevity. The $uu$ and $vv$ field equations, $G_{uu}=T_{uu}$ and $G_{vv}=T_{vv}$ yield constraints along the outgoing and ingoing null directions. [14]
For $uu$ component,
$r_{uu}^{{}^{\prime}}-\frac{f_{u}^{{}^{\prime}}}{f}r_{u}^{{}^{\prime}}=T_{uu}.$ (35)
For our matter, all $T_{uu}$ comes from outgoing null radiation as the EM field has no $T_{uu}$ in this direction, so $T_{uu}=\Psi(u)(l_{u})^{2}$. This equation can be rewritten as:
$\partial_{u}\Big{(}\frac{r_{u}^{{}^{\prime}}}{f}\Big{)}=-\frac{\Psi(u)(l_{u})^{2}}{f}$ (36)
which in vacuum integrates to $\frac{r_{u}^{{}^{\prime}}}{f}=A(v)$, with $A(v)$ an arbitrary function of $v$. Thus, in our case we have:
@@
and if the radiation flux is expressed in terms of the mass loss, we expect $T_{uu}=\Psi(u)\sim-\partial_{u}m(u)/(4\pi r^{2})$. In fact, integrating $G_{uu}$ reveals that $m(u)$ is an integration function related to the energy flux. In the pure radiation case ($q=0$), one indeed finds $r_{,u}=A(v)f$ (no explicit $u$-dependence aside from $f$) consistent with Vaidya metric in outgoing coordinates.
@@
$r_{vv}^{{}^{\prime}}-\frac{f_{v}^{{}^{\prime}}}{f}r_{v}^{{}^{\prime}}=T_{vv}.$ (38)
In our scenario with only outgoing radiation, we have no ingoing flux ($T_{vv}=0$). This integrates similarly to give $\frac{r_{v}^{{}^{\prime}}}{f}=B(u)$.
One can fix the scaling of $v$ such that $B(u)=1$. Physically, in the absence of ingoing matter, $r_{v}^{{}^{\prime}}$ can be adjusted by reparameterizing $v$. Below we will often choose this gauge so that $r_{v}^{{}^{\prime}}=f$ for simplicity.
@@
$\frac{\partial^{2}}{\partial u\partial v}(\ln f(u,v))+\frac{2\,r_{uv}^{{}^{\prime}}}{r}-\frac{2\,q(u)^{2}\,f(u,v)}{r^{4}}=0\,.$ (39)
@@
$\frac{\partial^{2}}{\partial u\partial v}(\ln f(u,v))+\frac{2r_{uv}^{{}^{\prime}}}{r}=0,$ (40)
@@ VI Conclusion
Overall, this work not only provides new insights into the theoretical underpinnings of radiating charged spacetimes but also lays the groundwork for future investigations. Further studies might explore extensions to rotating white holes, incorporate quantum gravitational effects, or examine the implications for astrophysical phenomena. The methodologies and results presented here contribute to a deeper understanding of how classical general relativity can be extended to accommodate dynamic processes such as evaporation, ultimately enriching our comprehension of extreme gravitational systems.

### v2

@@ III.2 Mass Loss and Outgoing Energy Flux with Hiscock Model
In the outgoing Vaidya picture, $M(u)$ encapsulates the evaporation law. For example, one might prescribe a linear decrease $M(u)=M_{0}-\alpha u$ during some interval or an asymptotic decay $M(u)\sim M_{0}-\beta\ln(1+u)$, etc. The rate $|M^{\prime}(u)|$ is related to the power carried by the radiation. In particular, far from the hole, the luminosity measured at infinity is $-\frac{dM}{du}$ (since $M^{\prime}(u)<0$). This matches the idea that the mass-energy radiated to infinity reduces the Bondi mass at future null infinity. The outgoing energy flux density is $T_{uu}=-M^{\prime}(u)/\left(4\pi r^{2}\right)$, integrating $4\pi r^{2}T_{uu}$ over a sphere at large $r$ reproduces the total luminosity $L(u)=-M^{\prime}(u)$, precisely the Bondi mass-loss rate.[13]
W. A. Hiscock (1981) [14] constructed explicit piece-wise Vaidya space-times to model an evaporating black hole. The geometry is glued together from three null regions.
Region I: an initial Schwarzschild space-time of mass $M_{0}$.
Region II: an ingoing Vaidya patch with a monotonically decreasing mass function $M(v)$ (advanced time), representing a stream of negative-energy null dust that encodes the Hawking back-reaction.
Region III: flat Minkowski space reached when $M(v)\rightarrow 0$.
This shown as figure 4.
Figure 3: Choice for $M(v)$ for the Hiscock model[14]
Because $M(v)$ vanishes at a finite advanced time $v_{\text{end }}$, the event horizon exists only for a finite interval and terminates at a point, beyond which a Cauchy horizon forms. Hiscock showed that for any evaporation law that brings the mass to zero in finite time-even a smooth linear decay-the renormalised stress tensor of a quantum test field diverges as
$\left\langle T_{ab}\right\rangle\propto\left(v_{\text{end }}-v\right)^{-2}$ (28)
on that Cauchy horizon, signaling an infinite burst of outgoing particles. The lesson is that a self-consistent semiclassical evaporation should have
$M(v)\longrightarrow 0\text{ only as }v\rightarrow+\infty,$ (29)
so that the horizon ”evaporates away” gradually without generating unbounded energy densities.
@@
At large $r$, $\theta_{(n)}\approx-1/r<0$ which ingoing rays converge as in flat space. But if we approach the radius $r=2M(u)$, the ingoing expansion $\theta_{(n)}$ approaches zero and changes sign. Inside the white hole, one actually has $\theta_{(n)}>0$ (an anti-trapped region where even ingoing light is forced outward). The surface $r=2M(u)$ thus acts as a marginal surface. In fact, $r=2M(u)$ is a dynamical apparent horizon for the spacetime: it is the boundary between normal space and the anti-trapped region. For the ingoing Vaidya black hole case, $r=2M(v)$ is a marginally trapped surface. In either case, the Vaidya horizon satisfies conditions like $\theta_{(\ell)}\theta_{(n)}=0$ on it. We identify $r_{h}(u)=2M(u)$ as the apparent horizon radius at a given retarded time $u$. Differentiating this:
@@ III.3 Horizon Dynamics and Spacetime Structure during Evaporation
In a dynamical space-time the event horizon is defined globally, whereas an apparent (or trapping) horizon is defined locally. For the outgoing-Vaidya white hole the natural apparent/anti-trapping horizon is the surface
$r=2M(u)$ (34)
which is a timelike past outer trapping horizon. If the evaporation leaves no remnant, the event horizon initially coincides with this surface but terminates when the mass function reaches zero. Concretely, if $M(u)$ vanishes at some finite retarded time $u_{\text{end }}$, then for $u>u_{\text{end }}$ the space-time is flat Minkowski and contains no trapped or anti-trapped surfaces; the event horizon therefore ends at that point.
If instead $M(u)$ tends to zero only asymptotically as $u\rightarrow\infty$, the apparent and event horizons shrink continuously to zero radius in the infinite future.
@@
In summary, for evaporating black hole, horizon is found to be timelike in the Vaidya solution (to first order in $M^{\prime}$). This means an outside observer could actually cross the retreating horizon if they move fast enough outward, since the horizon is moving inward. Intuitively, the horizon is contracting, so outgoing light just outside can outrun it. A timelike horizon also implies it has a surface area that decreases with time, when $dA<0$ we say it as timelike membrane. This is a violation of the classical area theorem, but it is possible here because the null energy condition and non-null horizon is violated by Hawking-like flux.
@@
For shrinking white hole, horizon is spacelike in analyses of dynamic solutions. In the time-forward picture, it means the horizon’s boundary moves at faster-than-light rate inward. In essence, the white hole’s anti-trapping horizon vanishes so quickly that it lies within the backward light cone of its boundary. Physically, once the white hole has emitted most of its mass, the remaining horizon collapses “faster than light” in the diagram, so no world line can remain on the horizon, as it is a fleeting, space-like boundary. This corresponds to a rapid evaporation or an endpoint of the white hole. By contrast, an expanding white hole, which would have increasing $M$ if such were considered, yields a timelike horizon, analogous to an evaporating Black hole’s horizon being timelike.
@@
These properties were confirmed in recent studies of Kerr–Vaidya spacetime[16], which reduce to spherical symmetry in the limit $a\to 0$, as $a$ defined as $a=J/M$ is the angular momentum per unit, as non-rotation circumstance, $a=0$. There, it was found that a shrinking white hole’s horizon is spacelike, whereas a slowly expanding one is timelike. Similar, a slowly evaporating black hole’s apparent horizon is timelike. All these dynamic horizons differ from the static case, which the classical event horizon will wrinkles and tilts due to the mass flux.
III.4 Global structure
@@
The spacetime of an evaporating white hole can be visualized as follows. Start with a past singularity (much like the white hole singularity in the past of an eternal Schwarzschild). Emanating from this $r=0$ singularity is the white hole event horizon (a past horizon) which initially might be near $r=2M_{\text{initial}}$. As time progresses, the horizon moves outward for a while (if the white hole initially increases to $M_{\text{initial}}$) and then begins to shrink as $M$ decreases. Outside the horizon, there is an outflow of radiation to infinity. If the evaporation is complete, at some retarded time $u=u_{+}$ the mass approaches zero. At this point, the horizon meets the center $r=0$. In a Penrose diagram, the horizon might terminate at a point on future null infinity or at future timelike infinity depending on how $M$ vanishes. In any case, after evaporation, the spacetime becomes flat or at least free of any horizon or singularity – effectively connecting to a region isometric to a portion of Minkowski space. No future singularity remains, since the original singularity was in the past. In this sense, the white hole has completely disappeared, leaving behind empty space.
@@ IV Vaidya-Reissner-Nordström white hole
We work in units $18\pi G=c=$ 1 and assume $r(u,v)$ and $f(u,v)$ are smooth. With $g_{uv}=-f$, the non-zero Einstein equations along the null directions are[18]
$G_{uu}=T_{uu},\quad G_{vv}=T_{vv}.$
The $uu$ component be:
$r_{uu}-\frac{f_{u}}{f}r_{u}=-\frac{f}{2r}T_{uu}.$ (37)
For our matter content the only $uu$-stress comes from outgoing null dust,
$T_{uu}=\Psi(u)\left(l_{u}\right)^{2}=\Psi(u),\quad l_{\mu}=-\partial_{\mu}u$ (38)
Dividing Eq.(37) by $f$ gives the differential form:
$\partial_{u}\left(\frac{r_{u}}{f}\right)=-\frac{\Psi(u)}{2r(u,v)}.$ (39)
Integrate Eq.(39) from a reference retarded time $u_{0}$ :
@@
where $A(v)$ is an arbitrary function of the advanced coordinate, fixed by the remaining gauge freedom or initial data.
Multiplying by $f(u,v)$ yields
[…]
In vacuum Eq.(40) reduces to $r_{u}/f=A(v)$, reproducing the familiar constraint for double null Schwarzschild.
If the radiation flux is expressed in terms of the mass loss, we expect $T_{uu}=\Psi(u)\sim-\partial_{u}m(u)/(4\pi r^{2})$ by assume $f=1$ In fact, integrating $G_{uu}$ reveals that $m(u)$ is an integration function related to the energy flux. In the pure radiation case ($q=0$), one indeed finds $r_{u}=A(v)f$ (no explicit $u$-dependence aside from $f$) consistent with Vaidya metric in outgoing coordinates.
@@
$r_{vv}-\frac{f_{v}}{f}r_{v}=-\frac{f}{2r}T_{vv}.$ (42)
In our scenario with only outgoing radiation, we have no ingoing flux ($T_{vv}=0$). This integrates similarly to give $\frac{r_{v}}{f}=B(u)$.
One can fix the scaling of $v$ such that $B(u)=1$. Physically, in the absence of ingoing matter, $r_{v}$ can be adjusted by reparameterizing $v$. Below we will often choose this gauge so that $r_{v}=f$ for simplicity.
@@
$\frac{\partial^{2}}{\partial u\partial v}(\ln f(u,v))+\frac{2r_{uv}}{r}-\frac{2q(u)^{2}f(u,v)}{r^{4}}=0.$ (43)
@@
$\frac{\partial^{2}}{\partial u\partial v}(\ln f(u,v))+\frac{2r_{uv}}{r}=0,$ (44)
@@ VI Conclusion
Overall, this work not only provides new insights into the theoretical underpinnings of radiating charged spacetime but also lays the groundwork for future investigations. Further studies might explore extensions to rotating white holes, incorporate quantum gravitational effects, or examine the implications for astrophysical phenomena. The methodologies and results presented here contribute to a deeper understanding of how classical general relativity can be extended to accommodate dynamic processes such as evaporation, ultimately enriching our comprehension of extreme gravitational systems.